# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ 3 GD with Armijo Line-search
We consider using gradient descent (GD) with Armijo line-search  (Armijo, 1966) (henceforth referred to as GD-LS) to minimize functions satisfying Assumptions 1, 2 and 3. The GD-LS update at iteration $t\in[T]$ is given as:
@@
where ${\eta_{t}}$ is the step-size returned by backtracking Armijo line-search (Armijo-LS). In particular, starting from an initial maximum step-size $\eta_{\max}$, Armijo-LS uses backtracking to select a step-size that satisfies the Armijo condition,
@@
where $c\in(0,1)$ is a tunable parameter. The complete pseudo-code is described in Algorithm 1. The parameter $\beta$ controls the backtracking and is typically set to $0.9$, while the parameter $c$ is typically set to a small value such as $10^{-4}$ (Nocedal & Wright, 2006).
@@
When using GD-LS for minimizing $L$ (uniformly)-smooth functions (corresponding to $L_{0}\neq 0$ and $L_{1}=0$ in Assumption 2), ${\eta_{t}}$ is constrained to lie in the $\left[\min\left\{\eta_{\max},\frac{2\,(1-c)\,\beta}{L}\right\},\eta_{\max}\right]$ range. Note that this bound holds for all $L$ (uniformly)-smooth functions, does not require convexity, and guarantees that backtracking line-search will terminate at a non-zero step-size. The parameter $c$ controls the “aggressiveness” of the algorithm; small $c$ values encourage a larger step-size. Hence, Armijo-LS can be seen as method to obtain a step-size proportional to $1/L$ without the knowledge of the global smoothness constant.
The bounds on the step-size can be used to derive the convergence rate for GD-LS. For example, for uniformly $L$-smooth and convex functions, the standard analysis shows that GD-LS converges to the optimum at an $O(1/T)$ rate (Nocedal & Wright, 2006), thus matching the rate of GD with a constant step-size equal to $1/L$ (henceforth referred to as GD(1/L)). However, as alluded to in Section 1, Armijo-LS enables GD to adapt to the “local” smoothness $L(\theta_{t})$ (the smoothness around iterate $\theta_{t}$), and results in faster convergence in practice. By studying non-uniform smooth functions satisfying Assumption 2, we aim to formally characterize this fast convergence.
To that end, we first show that when minimizing non-uniform smooth functions satisfying Assumption 2, Armijo-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. It is straightforward to relax this assumption similar to the standard analysis (Nocedal & Wright, 2006). We first prove the following lemma on the minimum step-size returned by Armijo-LS.
@@
If $f$ satisfies Assumptions 1, 2 and 3, at iteration $t$, GD-LS returns a step-size
@@
For “smaller” step-sizes $\eta\leq\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$, we use Eq. 1 to prove that the step-size will satisfy the Armijo condition, and the backtracking line-search will terminate. However, since Eq. 1 is only valid when $\left\|y-x\right\|$ is small, we cannot directly use it when the step-size returned by Armijo-LS is “large”. For this, we first prove that $g(\theta):=\ln(f(\theta))$ is $L_{1}$ uniformly-smooth. We then show that the step-size returned by the line-search on $f(\theta)$ satisfies an equivalent Armijo condition on $g(\theta)$, and use the uniform smoothness of $g(\theta)$ to lower-bound the step-size. ∎
Note that when $L_{1}=0$, the lower-bound on ${\eta_{t}}$ is similar to that for uniformly smooth functions. However, when $L_{0}=0$ (e.g. for logistic regression or multi-class classification), the lower-bound on ${\eta_{t}}$ is proportional to $1/f(\theta_{t})$, meaning that as $f(\theta_{t})$ decreases, the step-size returned by Armijo-LS increases. This enables the faster convergence of GD-LS.
We now present a meta-theorem (proved in Appendix B) which we will instantiate and interpret for convex losses in Section 4 and non-convex losses in Section 5.
@@
For a fixed $\epsilon>0$, if $f$ satisfies Assumptions 1, 2 and 3 and if for a constant $R>0$, $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\frac{[f(\theta)-f^{*}]^{2}}{R}$, ${f^{*}}>0$, then GD-LS with $\eta_{\max}=\infty$ requires
@@
$f(\theta_{t+1})\leq f(\theta_{t})-\frac{[f(\theta_{t})-f^{*}]^{2}}{[\lambda_{0}+\lambda_{1}\,f(\theta_{t})]\,R}$ (7)
We split the proof into two cases: Case (1) when $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$ for all $t\in[T]$ iterations required to obtain the desired sub-optimality. In this case, $\lambda_{1}\,f(\theta_{t})\geq\lambda_{0}$. Using this relation with Eq. 7 and following the arguments in Axiotis & Sviridenko (2023, Theorem 5.2) allows us to complete the proof of Case (1). For Case (2), we follow the proof of the first case for iterations $t\in[\tau]$ for which $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$, and obtain a similar rate. This corresponds to Phase 1. with faster convergence. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation with Eq. 7 and following the standard proof for uniformly smooth functions completes the proof for Phase 2. which results in slower convergence. Putting together the results for both phases completes the proof of Case (2). ∎
We note that the above theorem requires an additional condition which lower-bounds the gradient norm in terms of the function sub-optimality. In Section 4, we use convexity to satisfy this condition, whereas in Section 5, we use gradient domination to satisfy it. We also require $\eta_{\max}$ to be larger than $\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$ for all $t$. This ensures that the step-size is not constrained by the initial choice, but rather by the properties of the function. For conciseness, we express this condition as $\eta_{\max}=\infty$, and note that it is straightforward to relax it, albeit at the cost of clarity.
In order to interpret the above theorem, let us first consider the setting corresponding to $\lambda_{1}=0$. Here, case (2) is active, and the algorithm requires $O(1/\epsilon)$ iterations to achieve the desired sub-optimality, matches the standard result for uniformly smooth functions. Now consider the setting when $\lambda_{0}=0$. Here, case (1) is active, and GD-LS requires $O\left(R\,\left(\frac{f^{*}}{\epsilon}\right)\,\ln\left(\frac{1}{\epsilon}\right)\right)$ iterations. The iteration complexity thus depends on $f^{*}$, and in cases where $f^{*}$ is small, GD-LS can result in an improved rate. As a concrete example, consider the case when $f^{*}=\delta\,\epsilon$ where $\delta\geq 1$ is a constant independent of $\epsilon$. In this setting, GD-LS will result in a faster $O\left(R\,\ln\left(\frac{1}{\epsilon}\right)\right)$ convergence. Note that GD(1/L) does not benefit from such adaptivity, and will always result in a sublinear rate. For non-zero $\lambda_{0}$ and $\lambda_{1}$, the resulting rate depends on the value of $f^{*}$. If $f^{*}$ is larger than the threshold $\frac{2\lambda_{0}}{\lambda_{1}}$, GD-LS can result in the potentially fast rate corresponding to case (1) whereas if $f^{*}$ is smaller than the threshold, the algorithm has a two-phase behaviour: fast convergence until the loss becomes smaller than the threshold followed by slow convergence to the minimizer.
@@ 4 GD-LS for Convex Losses
In this section, we characterize the convergence rate of GD-LS on convex losses satisfying Assumptions 1, 2 and 3. Recall that binary classification using logistic regression and multi-class classification using the cross-entropy loss both satisfy Assumptions 1, 2 and 3 with $L_{0}=0$ and $\omega=0$. Consequently, we instantiate Theorem 1 for this setting and prove the following corollary in Appendix C.
@@
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex and satisfies Assumptions 1, 2 and 3 with $L_{0}=0$ and $\omega=0$, GD-LS with $\eta_{\max}=\infty$, requires $T\geq$
$\max\{2\lambda_{1}\,\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2},1\}\,\left(\frac{f^{*}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-f^{*}}{\epsilon}\right)$
iterations to ensure to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
@@
Referring to the explanation following Theorem 1, we conclude tht GD-LS can result in faster convergence over GD(1/L) when $f^{*}$ is small (see Fig. 1 for an experimental validation). In order to compare the result in Corollary 1 with existing works, let us consider the special case of logistic regression. In this case, GD-LS matches the rate for a variant of normalized gradient descent (NGD) in Axiotis & Sviridenko (2023, Theorem 5.2). However, unlike GD-LS, NGD requires the knowledge of $L_{1}$ making it relatively difficult and less likely to be implemented in practice. Furthermore, we note NGD is a specialized algorithm that is helpful to attain faster rates for certain problems (Mei et al., 2021; Hazan et al., 2015; Wilson et al., 2019), whereas GD-LS is universally used and can automatically exploit the problem structure. Moreover Corollary 1 is more general and also holds for the exponential loss.
While GD(1/L) can result in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov et al., 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 00Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
Consequently, we consider the special case when the data is linearly separable and $f^{*}=0$ and better characterize the fast convergence for logistic regression. Unfortunately, we cannot directly use Corollary 1 since $\left\|\theta\right\|\to\infty$ as $f(\theta)\to 0$, making the resulting bound vacuous (Orabona, 2024). Consequently, we use a different technique and first prove the following theorem in Appendix C.
@@
For a fixed $\epsilon>0$ and an arbitrary comparator $u$, if $f(\theta)$ is convex, $L$-uniform smooth and satisfies Assumptions 1, 2 and 3 with $L_{0}=0,\,\omega=0$, GD-LS with $\eta_{\max}=\infty$ and $c>\frac{1}{2}$, requires
@@
Compared to Corollary 1, the above result only focuses on the case when $L_{0}=0,\,\omega=0$ and requires an additional assumption that $f(\theta)$ is $L$-smooth. These conditions are satisfied for both logistic regression and multi-class classification. We use the above result and prove the following corollary for logistic regression on separable data.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon>0$, GD-LS with $\eta_{\max}=\infty$ requires
$T\geq\frac{6\,c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
Proof Sketch.
We use a proof technique similar to that in Ji & Telgarsky (2018), and consider the max-margin solution $u^{*}$ where $\left\|u^{*}\right\|=1$ and $\gamma=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$. Note that for any scalar $\beta>0$, $f(\beta u^{*})\leq\exp(-\beta\gamma)$. Choosing $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, setting the comparator $u=\beta\,u^{*}$ in Theorem 2 and using the bound on $f(\beta u^{*})$ completes the proof. ∎
Hence, on linearly separable data, GD-LS can achieve a linear rate of convergence for logistic regression (see Fig. 2 for an experimental validation), resulting in an exponential improvement over the standard $O(1/\epsilon)$ rate for smooth, convex functions. Finally, the result in Corollary 7 is better than the $O(1/\sqrt{\epsilon})$ rate for GD with large (than $1/L$) constant step-sizes (Wu et al., 2024).
Figure 2: Comparing GD-LS with $c=1/2$ and $\eta_{\max}=10^{8}$ and GD(1/L) for unregularized logistic regression on a synthetic separable dataset with $\gamma=0.1$, $n=10^{4}$ and $d=200$. (Left) Sub-optimality plot: GD-LS converges linearly, while GD(1/L) has a sublinear convergence. (Right) Step-size plot: The GD-LS step-size increases non-monotonically and becomes larger than $10^{5}$ in $40$ iterations.
An interesting question is whether the above linear convergence rate can only be achieved by methods such as GD-LS which ensure that the loss is monotonically decreasing. In Section C.1, we answer this in the negative, and show that the commonly used Polyak step-size (Polyak, 1987) can also achieve the fast convergence rate in Theorem 2 and consequently, result in linear convergence for logistic regression on separable data.
In the next section, we consider the convergence of GD-LS on non-convex losses.

### v2

@@ 3 GD with Armijo Line-search
The update for gradient descent (GD) with Armijo line-search  (Armijo, 1966) (henceforth referred to as GD-LS) at iteration $t\in[T]$ is given as:
@@
where ${\eta_{t}}$ is the step-size returned by the backtracking Armijo line-search (referred to as Armijo-LS). In particular, starting from an initial maximum step-size $\eta_{\max}$, Armijo-LS uses backtracking to select the (approximately) largest step-size that satisfies the Armijo condition,
@@
where $c\in(0,1)$ is a tunable parameter. The complete pseudo-code is described in Alg. 1. The parameter $\beta$ controls the backtracking and is typically set to $0.9$, while the parameter $c$ is typically set to a small value such as $10^{-4}$ (Nocedal & Wright, 2006).
@@
When using GD-LS for minimizing $L$ uniformly-smooth functions (corresponding to $L_{0}\neq 0$ and $L_{1}=0$ in Assn. 2), ${\eta_{t}}$ is constrained to lie in the $\left[\min\left\{\eta_{\max},\frac{2\,(1-c)\,\beta}{L}\right\},\eta_{\max}\right]$ range (Nocedal & Wright, 2006). Note that this bound holds for all $L$ uniformly-smooth functions, does not require convexity, and guarantees that the backtracking line-search will terminate at a non-zero step-size. The parameter $c$ controls the “aggressiveness” of the algorithm; small $c$ values encourage a larger step-size. Hence, Armijo-LS can be seen as method to obtain a step-size proportional to $1/L$, without the knowledge of the global smoothness constant.
These bounds on the step-size can be used to derive the convergence rate for GD-LS. For example, for uniformly $L$-smooth and convex functions, the standard analysis shows that GD-LS converges to an optimum at an $O(1/T)$ rate (Nocedal & Wright, 2006). It thus matches the rate of GD with a constant step-size equal to $1/L$ (henceforth referred to as GD(1/L)). However, as alluded to in Sec. 1, Armijo-LS enables GD to adapt to the “local” smoothness $L(\theta_{t})$ (the smoothness around iterate $\theta_{t}$), and results in faster convergence both in theory (Scheinberg et al., 2014; Lu & Mei, 2023; Mishkin et al., 2024; Fox & Schmidt, ), and in practice. In contrast to these works that show constant factor improvements for GD-LS, we study non-uniform smooth functions satisfying Assn. 2 and 3, and aim to show that GD-LS can result in a faster rate of convergence.
To that end, we first show that, when minimizing non-uniform smooth functions satisfying Assn. 2 and 3, GD-LS can result in provably larger step-sizes that enable faster convergence. For the subsequent theoretical analysis, we only consider “exact backtracking” i.e. we assume that the backtracking procedure returns the largest step-size that satisfies the Armijo condition, meaning that $\beta\approx 1$. Similar to the standard analysis (Nocedal & Wright, 2006), it is straightforward to relax this assumption
We first prove the following lemma that lower-bounds the step-size returned by GD-LS.
@@
If $f$ satisfies Assn. 1, 2 and 3, at iteration $t$, GD-LS returns a step-size
@@
For “smaller” step-sizes $\eta\leq\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$, we use Eq. 1 to prove that the step-size will satisfy the Armijo condition, and the backtracking line-search will terminate. However, since Eq. 1 is only valid when $\left\|y-x\right\|$ is small, we cannot directly use it when the step-size returned by Armijo-LS is “large”. Consequently, we first prove that $g(\theta):=\ln(L_{0}+L_{1}\,f(\theta))$ is $L_{1}$ uniformly-smooth. We then show that the step-size returned by the line-search on $f(\theta)$ satisfies an equivalent Armijo condition on $g(\theta)$. Using the uniform smoothness of $g(\theta)$ along with the standard line-search analysis enables us to lower-bound the step-size. ∎
Note that when $L_{1}=0$, the lower-bound on ${\eta_{t}}$ is similar to that for uniformly smooth functions. However, when $L_{0}=0$ and $\omega=0$ (e.g. for logistic regression in Prop. 1), the lower-bound on ${\eta_{t}}$ is proportional to $1/f(\theta_{t})$, meaning that as $f(\theta_{t})$ decreases, the step-size returned by Armijo-LS increases. This enables the faster convergence of GD-LS.
We now present a meta-theorem (proved in App. B) that characterizes this fast convergence for both convex and non-convex functions. It requires an additional condition which lower-bounds the gradient norm in terms of the function sub-optimality. In Sec. 4, we use convexity to satisfy this condition, whereas in Sec. 5, we use gradient domination to satisfy it. We also require that the step-size is not constrained by the initial choice, but rather by the properties of the function. This can be achieved by using a large $\eta_{\max}$, which is greater than $\frac{1}{\lambda_{0}+\lambda_{1}\,f(\theta_{t})}$ for all $t$, or by using a forward-tracking line-search to (approximately) return the largest step-size satisfying the Armijo condition (Fridovich-Keil & Recht, 2019). For conciseness, we express this requirement as $\eta_{\max}=\infty$. We note that it is straightforward to relax it and get an explicit dependence on $\eta_{\max}$, albeit at the cost of clarity in our theoretical results.
@@
For a fixed $\epsilon>0$, if $f$ satisfies Assn. 1, 2 and 3, and if for a constant $R>0$, $\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\geq\frac{[f(\theta_{t})-f^{*}]^{2}}{R}$ for all iterations $t\in[T]$, then, GD-LS with $\eta_{\max}=\infty$ requires
@@
$f(\theta_{t+1})\leq f(\theta_{t})-\frac{[f(\theta_{t})-f^{*}]^{2}}{[\lambda_{0}+\lambda_{1}\,f(\theta_{t})]\,R}.$ (7)
We now split the proof into two cases: Case (1) when $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$ for all $t\in[T]$. In this case, $\lambda_{1}\,f(\theta_{t})\geq\lambda_{0}$. Using this relation with Eq. 7 and following the arguments in Axiotis & Sviridenko (2023, Theorem 5.2) allows us to complete the proof of Case (1). For Case (2), we follow the proof of the first case for iterations $t\in[\tau]$ for which $f(\theta_{t})\geq\frac{\lambda_{0}}{\lambda_{1}}$, and obtain a similar rate. This corresponds to Phase 1. with faster convergence. For all iterations $t\in[\tau,T]$, $f(\theta_{t})\leq\frac{\lambda_{0}}{\lambda_{1}}$ and hence $\lambda_{0}\geq\lambda_{1}\,f(\theta_{t})$. Using this relation with Eq. 7 and following the standard proof for uniformly smooth functions completes the proof for Phase 2. which results in slower convergence. Putting the results for both phases together completes the proof of Case (2). ∎
In order to interpret the above theorem, let us first consider the setting corresponding to $\lambda_{1}=0$. Here, case (2) is active, and the algorithm requires $O(1/\epsilon)$ iterations to achieve the desired sub-optimality, matching the standard result for uniformly smooth functions. Now consider the setting when $\lambda_{0}=0$. Here, case (1) is active, and GD-LS requires $O\left(R\,\left(\frac{f^{*}}{\epsilon}\right)\,\ln\left(\frac{1}{\epsilon}\right)\right)$ iterations. The iteration complexity thus depends on $f^{*}$, and in cases where $f^{*}$ is small, GD-LS can result in an improved rate. As a concrete example, consider the case when $\epsilon=\Theta(f^{*})$. In this setting, GD-LS will result in a faster $O\left(R\,\ln\left(\frac{1}{\epsilon}\right)\right)$ convergence. Note that GD(1/L) does not benefit from such adaptivity, and will always result in a sublinear rate. For non-zero $\lambda_{0}$ and $\lambda_{1}$, the resulting rate depends on the value of $f^{*}$. If $f^{*}$ is larger than the threshold $\frac{2\lambda_{0}}{\lambda_{1}}$, GD-LS can result in the potentially fast rate corresponding to case (1), whereas if $f^{*}$ is smaller than the threshold, the algorithm has a two-phase behaviour: fast convergence until the loss becomes smaller than the threshold, followed by slow convergence to the minimizer.
@@ 4 GD-LS for Convex Losses
In this section, we characterize the convergence rate of GD-LS on convex losses satisfying Assn. 1, 2 and 3. Recall that binary classification using logistic regression and multi-class classification using the cross-entropy loss both satisfy Assn. 1, 2 and 3 with $L_{0}=0$ and $\omega=0$. Consequently, we only instantiate Thm. 1 for this setting and prove the following corollary in App. C.
@@
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex and satisfies Assn. 1, 2 and 3 with $L_{0}=0$ and $\omega=0$, GD-LS with $\eta_{\max}=\infty$, requires
$T\geq\max\{2\lambda_{1}\,\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2},1\}\,\left(\frac{f^{*}}{\epsilon}+1\right)\,\ln\left(\frac{f(\theta_{0})-f^{*}}{\epsilon}\right)$
iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
@@
Referring to the explanation following Thm. 1, we conclude tht GD-LS can result in faster convergence than GD(1/L) when $f^{*}$ is small (see Fig. 1 for an experimental validation).
In order to compare the result in Cor. 1 with existing works, consider the special case of logistic regression. In this case, GD-LS matches the rate for a variant of normalized gradient descent (NGD) in Axiotis & Sviridenko (2023, Theorem 5.2). However, unlike GD-LS, NGD requires the knowledge of $L_{1}$ making it relatively difficult to be implemented in practice. Furthermore, we note NGD is a specialized algorithm that is helpful to attain faster rates for certain problems (Mei et al., 2021; Hazan et al., 2015; Wilson et al., 2019), whereas GD-LS is universally used and can automatically exploit the problem structure. Moreover, Cor. 1 is more general and also holds for the exponential loss.
While GD(1/L) results in an $\Omega(1/\epsilon)$ rate for general smooth, convex functions (Nesterov et al., 2018), analyses of GD on logistic regression often exploit strong-convexity and prove faster rates (Karimi et al., 2016). In particular, if the iterates lie in a bounded set, the objective is $\mu(\theta)$ strongly-convex where $\mu(\theta)=\lambda_{\min}[X^{T}X]\,\min_{i}\pi_{i}(\theta)\,(1-\pi_{i}(\theta))$ where $\pi_{i}=\frac{1}{1+\exp(-y_{i}\,\langle x_{i},\theta\rangle)}$. Notice that as $\pi_{i}(\theta)$ tends to either zero or one i.e. the predictions become deterministic, $\mu(\theta)$ tends to 00Freund et al. (2018, Theorem 3.3) characterize the resulting rate for GD(1/L) and prove that the suboptimality scales as $O\left(\exp(-T\,\exp\left(\nicefrac{{-1}}{{\xi}}\right))\right)$ where $\xi$ is the degree of non-separability and tends to zero as the data becomes more separable. Hence, the rate becomes exponentially worse as $\xi$ decreases. However, as the data becomes separable, GD-LS converges at a faster linear rate (see Fig. 2), meaning that strong-convexity cannot explain this behaviour.
Consequently, we consider the special case where the data is linearly separable and $f^{*}=0$, and better characterize the fast convergence of GD-LS. Unfortunately, we cannot directly use Cor. 1 since $\left\|\theta\right\|\to\infty$ as $f(\theta)\to 0$, making the resulting bound vacuous (Orabona, 2024). Consequently, we use a different technique, and first prove the following theorem in App. C.
@@
For an initialization $\theta_{0}$, $\epsilon\in(0,f(\theta_{0}))$ and comparator $u$ s.t. $f(u)\leq\epsilon$, if $f(\theta)$ is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$, then, GD-LS with $\eta_{\max}=\infty$, $c>\frac{1}{2}$, requires
@@
Compared to Cor. 1, the above result only holds for a restricted range of $\epsilon$ and a comparator $u$ s.t. $f(u)\leq\epsilon$. Note that since $f$ is non-negative and GD-LS ensures monotonic descent, its sub-optimality is upper-bounded by $f(\theta_{0})$. The result only requires the non-uniform smoothness assumptions and thus holds for the exponential loss, logistic regression and multi-class classification.
We now use the above result and prove the following corollary for logistic regression on separable data (similar results hold for the exponential loss and multi-class classification).
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, an $\epsilon\in(0,f(\theta_{0}))$, GD-LS with $\eta_{\max}=\infty$ requires
$T\geq O\left(\frac{c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}\right)$
iterations to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
Hence, on linearly separable data, GD-LS can achieve a linear rate of convergence for logistic regression (see Fig. 2 for an experimental validation). In this setting, Wu et al. (2024, Theorem 3) show that GD(1/L) (or more generally, GD with any constant step-size that guarantees monotonic descent) cannot have a convergence rate faster than $\Omega(1/\epsilon)$. Hence, GD-LS can result in an exponential improvement over the rate for GD(1/L). Furthermore, the rate in Cor. 2 is better than the $O(1/\sqrt{\epsilon})$ rate for GD with large (beyond $1/L$) constant step-sizes (Wu et al., 2024). Finally, we note that, in this setting, no algorithm that guarantees monotonic descent in the function values can achieve a rate faster than the linear rate (Zhang et al., 2025, Theorem 2.2), Hence, in this sense, GD-LS is optimal for logistic regression.
Figure 2: Comparing GD-LS with $c=1/2$, $\eta_{\max}=10^{8}$ and GD(1/L) for unregularized logistic regression on a synthetic separable dataset with $\gamma=0.1$, $n=10^{4}$ and $d=200$. (Left) Sub-optimality plot: GD-LS converges linearly, while GD(1/L) has a sublinear convergence. (Right) Step-size plot: The GD-LS step-size increases non-monotonically.
Interestingly, in Sec. C.1, we prove that GD with the Polyak step-size (Polyak, 1987) (which does not necessarily result in monotonic descent in the function values) can also achieve the linear convergence rate in Thm. 2.
Convergence under $(L_{c},L_{g})$ non-uniform smoothness: Using the reduction in Prop. 3 allows characterizing the convergence of GD-LS on twice-differentiable, non-negative, $(L_{c},L_{g})$ non-uniform smooth (Zhang et al., 2019) and convex functions. In particular, by using Prop. 3 in conjunction with Thm. 1 allows us bound the sub-optimality for the last-iterate of GD-LS (see Cor. 8 in App. C for the formal statement) without requiring the knowledge of $L_{c}$ or $L_{g}$. In particular, assuming $L_{g}\geq 1$, $L_{c}\geq 1$, we prove that GD-LS requires $\tilde{O}\left(\frac{L_{c}\,L_{g}^{2}\,R}{\epsilon}+\frac{L_{g}\,(L_{g}^{2}+L_{c})\,R\,f^{*}}{\epsilon}\right)$ iterations. In contrast, in this same setting, Vankov et al. (2024); Gorbunov et al. (2024) show that GD with the Polyak step-size requires $O\left(\max\left\{\frac{L_{c}}{\epsilon},L_{g}\,R\right\}\right)$ iterations for the best-iterate to achieve an $\epsilon$ sub-optimality (Gorbunov et al., 2024, Thm. 4.1). Comparing the two results, we note that, in general, the upper-bound in (Gorbunov et al., 2024) is tighter in terms of $\epsilon$. However, when $\epsilon=\Theta(f^{*})$, the two bounds are equivalent in their $\epsilon$ dependence. Moreover, while the results in Gorbunov et al. (2024); Vankov et al. (2024) hold for the best-iterate of GD with the Polyak step-size (that requires the knowledge of $f^{*}$), Cor. 8 holds for the last-iterate and does not require knowing $f^{*}$.
Next, we analyze GD-LS on non-convex losses.